OpenStudy (anonymous):

Solve for x: |2x + 6| − 4 = 20

OpenStudy (anonymous):

First you need to isolate the absolute value part of the equation. So what would you do?

OpenStudy (anonymous):

I don't know. That's why i posted on here.

OpenStudy (anonymous):

What do you get if you add 4 to both sides of the equation?

OpenStudy (anonymous):

|2x + 6| − 4 = 20 |2x + 6| = 20 + 4 (|2x + 6|)^2= 24^2 2x+6 = + - sqrt(24^2) 2x +6 = + - 24 So 2x +6 = 24 2x = 24 - 6 2x = 22 x = 11 or 2x +6 = - 24 2x =- 24 - 6 2x = -30 x = -30/2 x= -15

OpenStudy (anonymous):

they are x = 11 and x = -15

OpenStudy (anonymous):

Solve for x: |x| − 8 = −5 what about this one?

OpenStudy (anonymous):

i cannot see it clearly, would you mind writing it again ?

OpenStudy (anonymous):

x -8 =-5

OpenStudy (anonymous):

x -8 =-5 adding both side by 8, then x -8 + 8 =-5 + 8 x = 3

OpenStudy (anonymous):

Thank you, thank you. Do you mind helping me on more?

OpenStudy (anonymous):

x=3 is only part of the answer.

OpenStudy (anonymous):

In solving absolute values take the equation and solve for both the value given and the value given times negative.....|x| - 8 = -5 to find both values for x you would find x-8= -5 and x-8 = 5.

OpenStudy (anonymous):

you are welcome.., go ahead @AnnastasiaElizabeth3 :)

OpenStudy (anonymous):

Solve for x: -3|x + 7| = -12

OpenStudy (anonymous):

@nelsonjedi when i told @AnnastasiaElizabeth3 "i cannot see it clearly, would you mind writing it again ?". She wrote x -8 =-5, not |x| -8 = -5

OpenStudy (anonymous):

-3|x+7| = -12 |x+7| = 12/3 = 4 (|x+7|)^2 = 4^2 x+7 = + - sqrt(4^2) x + 7 = + - 4 so x + 7 = 4 x = 4-7 x = -3 or x + 7 = - 4 x = -4 -7 x = -11 so the answer are x = -11 and x = -3

OpenStudy (anonymous):

get it @AnnastasiaElizabeth3 ?

OpenStudy (anonymous):

no. But I have one more for ya

OpenStudy (anonymous):

why no??

OpenStudy (anonymous):

he minimum and maximum temperature on a cold day in Lollypop Town can be modeled by the equation below: 2|x − 6| + 14 = 38 What are the minimum and maximum temperatures for this day?

OpenStudy (anonymous):

i cannot see it clearly again..,

OpenStudy (anonymous):

2 x-6 + 14=38

OpenStudy (anonymous):

i think it is 2 |x-6 |+ 14=38, right ?

OpenStudy (anonymous):

Yessss.

OpenStudy (anonymous):

2 |x-6| + 14 = 38 2|x-6| = 38 - 14 2|x-6| = 24 |x-6| = 24/2 |x-6| = 12 SO x - 6 = 12 x = 12 + 6 x = 18 OR x - 6 = -12 x = -12 + 6 x = -6

OpenStudy (anonymous):

the are x = 18 and x = -6

OpenStudy (anonymous):

THANK YOU SO MUCH!

OpenStudy (anonymous):

you are very welcome !!

OpenStudy (anonymous):

I have more but I don't want to bother you.

OpenStudy (anonymous):

you should bother me more with your quastion, go ahead., :p

OpenStudy (anonymous):

Solve for x: 3x - 5 = 2x + 6

OpenStudy (anonymous):

sorry for the bother.

OpenStudy (anonymous):

3x - 5 = 2x + 6 3x - 2x = 6 + 5 x = 11

OpenStudy (anonymous):

nevermind @AnnastasiaElizabeth3 :),

OpenStudy (anonymous):

i want to see you photo

OpenStudy (anonymous):

Solve for x: 3(x + 1) = -2(x - 1) - 4.

OpenStudy (anonymous):

3(x + 1) = -2(x - 1) - 4 3x + 3 = -2x + 2 - 4 3x + 2x = 2 - 4 - 3 5x = -5 x = -1

OpenStudy (anonymous):

@AnnastasiaElizabeth3 i've solved that

OpenStudy (anonymous):

Solve for x: -3(x + 3) = -3(x + 1) - 5.

OpenStudy (anonymous):

i am tired :(

OpenStudy (anonymous):

i am getting sleepy right now

OpenStudy (anonymous):

i'll solve it next time